# Massive Black Hole Seeds

### v1

@@ 1 Introduction
More recently, observations with JWST have uncovered MBHs with masses in the range $10^{6}$ $\rm{M_{\odot}}~{}$to $10^{8}$ $\rm{M_{\odot}}~{}$(Kovacs et al., 2024; Maiolino et al., 2024b; Larson et al., 2023; Bogdán et al., 2023; Matthee et al., 2023; Greene et al., 2024) at even higher redshifts. These MBH masses are lower than the high-z quasar masses previously detected but the galaxies within which these “JWST-quasars” have been detected likely represent the tip of the iceberg at this time in terms of both halo and stellar masses (Boylan-Kolchin, 2023; Juodžbalis et al., 2024; Li et al., 2024). Nonetheless, with inferred masses down as low as $5\times 10^{6}$ $\rm{M_{\odot}}~{}$(Maiolino et al., 2024b) these observations set an upper limit to the seed masses of this population of MBHs at the very least. The number density of these MBH host galaxies is currently unclear with potentially significant evolution of the central black hole following seeding. If we assume that current candidate AGN being observed by JWST are indeed true AGN then the number densities could be of the order $10^{-4}$ to $10^{-3}$ cMpc-3 (Matthee et al., 2023; Greene et al., 2024; Scholtz et al., 2023).The goal of this paper is to highlight what the current state-of-the-art in MBH seeding is, to identify challenges in the current models and to plot a path forward in terms of reconciling current theoretical models with observational data.
@@
Alternative routes to MBH formation may be through either a dynamical runaway pathway where either stellar collisions and/or black hole mergers result in a heavy seed. Another possibility is a near monolithic collapse, or at least reduced fragmentation of a massive gas cloud, resulting in the formation of a supermassive star (SMS). Both of these mechanisms create a heavy seed with masses in excess of approximately $10^{3}$ $\rm{M_{\odot}}~{}$and in some cases possibly up to $10^{5}$ $\rm{M_{\odot}}$. We will address these questions and gauge their possible number density in §3. We begin by first briefly reviewing the light seed model and then the heavy seed model. More in-depth reviews of the mechanisms can be found in reviews by Latif & Ferrara (2016); Woods et al. (2017); Inayoshi, Visbal & Haiman (2020); Volonteri, Habouzit & Colpi (2021). Additional mechanisms, which we will not consider in this paper, include supermassive stars fuelled by dark matter annihilation (e.g., Spolyar, Freese & Gondolo, 2008; Singh, Monaco & Tan, 2023) or primordial black holes (e.g., Carr & Silk, 2018; Hasinger, 2020).
@@ 2 Light Seeds as the origin of all MBHs
where $\rm{t_{Edd}}=0.45$ Gyr and $\epsilon_{r}$ is the radiative efficiency. For a “standard” radiative efficiency of $\epsilon_{r}\sim 0.1$ and a black hole seed mass of $\rm{M_{0}}=10^{2}$ $\rm{M_{\odot}}$, it takes nearly 1 Gyr to grow from $\rm{M(0)}$ up to $\sim 10^{9}$ $\rm{M_{\odot}}$. While therefore in principle marginally allowable with current observational constraints the growth must continue at the Eddington rate all the way up to the maximum mass (i.e. the duty cycle must be unity). Such a scenario has been shown to be exceedingly unlikely with numerous models and cosmological simulations showing that light seeds do not grow efficiently and tend to remain at their seed masses across cosmic time (Alvarez, Wise & Abel, 2009; Milosavljević, Couch & Bromm, 2009; Smith et al., 2018; Spinoso et al., 2022).
@@ 4 Seed Number Densities
Dijkstra, Ferrara & Mesinger (2014) calculated the number density of heavy seeds forming through the so-called Lyman-Werner (LW) channel. In this case an atomic halo is irradiated with a super-critical LW by a nearby galaxy (Dijkstra et al., 2008; Agarwal et al., 2012, 2013; Visbal, Haiman & Bryan, 2014; Agarwal et al., 2015; Regan et al., 2017). In principle this can lead to the formation of an extremely massive seed (see §5). However, the calculated number densities are small with expected number densities in the range $10^{-10}$ cMpc-3 to $10^{-5}$ cMpc-3 (at $z\sim 10$). It should be noted here that the higher number densities quoted here (i.e. $10^{-5}$ cMpc-3) should be treated as strong upper limits as they rely on very optimistic formation criteria, namely, low values of the LW radiation needed to keep $\rm{H_{2}}$ dissociated or allowing for metal pollution. Other authors have drawn similar conclusions on the expected number densities from this channel (Inayoshi & Tanaka, 2015; Habouzit et al., 2016; Agarwal et al., 2016; Trinca et al., 2022).
Moving to channels which potentially produce heavy seeds below the theoretical maximum mass (of SMSs) but well in excess of light seeds, the so-called rapid assembly channel (Yoshida et al., 2003; Fernandez et al., 2014; Wise et al., 2019; Regan et al., 2020c; Lupi, Haiman & Volonteri, 2021; Latif et al., 2022) can produce heavy seeds with number densities in the range 10-5 to $10^{-1}$ cMpc-3 (at $z\sim 10$) (McCaffrey et al. in prep) depending on environmental factors. Number densities achieved through other channels like baryonic streaming velocities result in number densities close to approximately 10-5 cMpc-3 (at $z\sim 10$) as well (Hirano et al., 2017).
@@ 5.1 The 1000J Halo
The number density of haloes subjected to a super-critical LW flux cannot be obtained from the simulations discussed above due to their idealistic nature. However, both Visbal, Haiman & Bryan (2014) and Regan et al. (2020b) (using the Renaissance simulation suite) computed the abundance of so-called synchronised pairs which can lead to the emergence of haloes similar to the 1000J halo (Regan et al., 2017). Both authors found the abundance of synchronised haloes to be less than $10^{-2}$ cMpc-3 (at z $\sim 10$). However, this number should be seen as an upper limit. For the "Normal" region of the Renaissance simulation suite Regan et al. (2020b) found exactly zero synchronised pairs while finding five, in the clustered, "Rarepeak" region. Taking into account the relative rarity of the Rarepeak region, the number densities obtained should be decreased by a factor of approximately 1000. This leads to number densities of less than $10^{-5}$ cMpc-3 (at z $\sim 10$) for a 1000J like object.
@@ 6 Model Halo Comparison
There are two primary reasons for the elevated stellar masses found in the 1000J compared to the Rapid and PopIII halo. On the one hand and as described above the LW flux dissociates $\rm{H_{2}}$ and leads to larger Jeans masses and hence larger protostars (Prole et al., 2023) but also these protostars can accrete more readily and for longer (see Figure 5). In the Rapid halo and PopIII halo the structures are far more asymmetric making continued accretion much more difficult (at least for the stellar masses probed here). For the cases of stars in both the Rapid and PopIII haloes the stars can initially accrete gas readily but the gas supply quickly gets consumed and the stars contract to the main sequence (Regan et al., 2020c) with masses significantly below that seen in more symmetric and homogeneous halos.
It should be noted that in Figure 5 we show only accretion over a relatively short period (hundreds of kilo-years). The subsequent growth of the seed (light or heavy) will depend fundamentally on the evolution of the host galaxy over several tens and hundreds of megayears. Presently, simulations able to track the model the formation of light and heavy seeds and to subsequently follow their growth are extremely computationally challenging and essentially intractable. We nonetheless describe progress in this direction in the following section bearing in mind that this remains an extremely challenging modelling problem.
Figure 5: The evolution of the most massive star that forms in each of the three haloes (1000J, Rapid and PopIII). In the top panel we show the mass as a function of time. In the bottom panel we show the accretion rate onto the most massive star in each halo as a function of time. The 1000J halo forms the most massive object followed by the Rapid halo and the PopIII halo. The dashed line shows the critical accretion that must be maintained onto the stellar surface to meet the criteria for being a super-massive star (Nandal et al., 2023).
@@ 7.1 Growth of Light Seeds
Smith et al. (2018) et al. used the Renaissance suite to track, in post-processing, the evolution of light seed black holes formed from PopIII stars. They tracked the evolution of over 20,000 light seeds for up to 200 Myrs and found negligible growth across the entire light seed population. The reason being is that the light seeds experience extremely low duty cycles due to the surrounding (negative) feedback processes of star formation and stellar feedback. Moreover, higher resolution studies of (PopIII) star forming environments show that PopIII stars form in multiples (e.g. Turk, Abel & O’Shea, 2009; Clark et al., 2011; Prole et al., 2023) setting off a competitive accretion scenario further mitigating against efficient accretion. Other authors (e.g. Pfister et al., 2019; Ma et al., 2021) find that only seeds with masses in the heavy seed (up to $10^{8}$ $\rm{M_{\odot}}~{}$in the case of the Ma et al. (2021) study) range can effectively sink to the centre of the potential well where efficient accretion may be sustainable. None of the above studies rule out light seed accretion but instead find that the probability of efficient light seed growth is extremely low.
Building on these challenging conclusions from studies of light seed growth a number of groups have investigated how growth can be accelerated when the conditions are favourable. In the majority of these cases the light seed black hole is modelled to either sit at the centre of a converging flow (e.g. Volonteri & Rees, 2005) or is "trapped" within the confines of a nuclear star cluster (Alexander & Natarajan, 2014; Natarajan, 2020) or within a circumnuclear disk (e.g. Lupi et al., 2016). More recent work by Shi, Kremer & Hopkins (2024) shows that a small fraction ($\sim$ 1%) of light seeds can achieve bursts of rapid growth when placed in environments that are favourable (i.e. are gas rich and support growth). In each of these cases the growth of the initially light seed proceeds via bursts of super-Eddington accretion. The physical process of super-Eddington accretion has been demonstrated extensively via extremely high resolution simulations (e.g. Sa̧dowski, 2009; Sa̧dowski & Narayan, 2016; Sa̧dowski et al., 2016; Jiang, Stone & Davis, 2017; Jiang et al., 2020). However, what is lacking are robust numerical simulations which show self-consistently the growth of light seeds in a cosmological setting. The reason for this is that a combination of high resolution, with resolved dynamics, and a realisation containing a rare dynamical occurrence is required. This has, thus far, not been realised. However, such a setup should be tractable with current simulation codes and techniques at least for the early evolutionary phase.
@@ 7.2 Heavy Seed Growth
So overall where does the extensive work on black hole growth leave us? Across the various simulation setups there is convergence that black holes can grow once the host halo becomes massive enough with some differences arising for black hole growth in lower mass galaxies which needs to be fully understood. Additionally, the black hole must achieve a mass of at least $10^{5}$ $\rm{M_{\odot}}~{}$before growth can be achieved - this is particularly evident from the Astrid simulations which have the lowest seed masses ($\rm{M_{seed}\gtrsim 10^{4}}$ $\rm{M_{\odot}}$) and see very limited growth for the lightest seeds. Similar results are obtained in semi-analytical models (Trinca et al., 2022; Spinoso et al., 2022). The main result therefore is that any MBH seeds below approximately $10^{5}$ $\rm{M_{\odot}}~{}$grow extremely inefficiently due to dynamics. In addition to this, in low mass galaxies ($\rm{M_{*}\lesssim 10^{9}}$ $\rm{M_{\odot}}~{}$) growth may by hampered by supernova feedback but this result suffers from different interpretations in the literature and requires further study.
@@ 8 Summary and Conclusions
The goal of this paper is to weave together recent developments in modeling MBH seeding and pathways to their growth. The current paradigm is broken into light and heavy seeds with light seeds typically defined as those with initial masses less than 1000 $\rm{M_{\odot}}~{}$and heavy seeds as those with initial masses in excess of 1000 $\rm{M_{\odot}}$. Scenarios considered in the past gave rise to a bimodal initial mass function: common light seeds related to PopIII stars with mass $\sim 100$ $\rm{M_{\odot}}~{}$and rarer heavy seeds related to SMSs with mass $\sim 10^{5}$ $\rm{M_{\odot}}$. In order to grow these seeds into the massive and supermassive regime, two main pathways were considered: super-Eddington on light seeds and sub-Eddington on heavy seeds, although of course super-Eddington accretion can occur, and indeed it may be more likely to occur, for heavy seeds. What the most recent simulations are instead suggesting is that these two regimes are likely to be the extreme cases of a broad distribution of initial properties.
@@
    Light Seed: Smaller haloes or those enriched with metals will produce relatively less massive stars. These stars will be distributed inhomogeneously around the halo centre and will accrete very inefficiently. The largest stellar masses shown here have masses up to a few times $\sim 10^{2}$ $\rm{M_{\odot}}~{}$(PopIII halo). The number densities of these halos are between $10^{-1}$ cMpc-3 and up to $10$ cMpc-3 (z $\sim$ 10).
@@
    Need for gas replenishment: In all three halos considered the accretion rate on the stellar progenitors of the seeds drops after $\sim 10^{5}$ years as the readily accretable gas in the parent halo is consumed. Gas replenishment is needed (Johnson & Bromm, 2007; Lupi et al., 2023) through either mergers with gas-rich halos or accretion from the cosmic web.
@@
    Heavy Seed Accretion: Heavy seeds experience the same bottlenecks as light seeds, but their effect is mitigated in relation to the seed masses: Bondi accretion scales as $M_{BH}^{2}$ and gravity scales as $M_{BH}$. A 10-times more massive seed has better chances of remaining in dense regions and accreting from them (Pfister et al., 2019; Lupi et al., 2023).

### v2

@@ 1 Introduction
More recently, observations with JWST have uncovered MBHs with masses in the range $10^{6}$ $\rm{M_{\odot}}~{}$to $10^{8}$ $\rm{M_{\odot}}~{}$(Kovács et al., 2024; Maiolino et al., 2024b; Larson et al., 2023; Bogdán et al., 2023; Matthee et al., 2024; Greene et al., 2024) at even higher redshifts. These MBH masses are lower than the high-z quasar masses previously detected but the galaxies within which these “JWST-quasars” have been detected likely represent the tip of the iceberg at this time in terms of both halo and stellar masses (Boylan-Kolchin, 2023; Juodžbalis et al., 2024; Li et al., 2024). Nonetheless, with inferred masses down as low as $5\times 10^{6}$ $\rm{M_{\odot}}~{}$(Maiolino et al., 2024b) these observations set an upper limit to the seed masses of this population of MBHs at the very least. The number density of these MBH host galaxies is currently unclear with potentially significant evolution of the central black hole following seeding. If we assume that current candidate AGN being observed by JWST are indeed true AGN then the number densities could be of the order $10^{-4}$ to $10^{-3}$ cMpc-3 (Matthee et al., 2024; Greene et al., 2024; Scholtz et al., 2023).The goal of this paper is to highlight what the current state-of-the-art in MBH seeding is, to identify challenges in the current models and to plot a path forward in terms of reconciling current theoretical models with observational data.
@@
Alternative routes to MBH formation may be through either a dynamical runaway pathway where either stellar collisions and/or black hole mergers result in a heavy seed. Another possibility is a near monolithic collapse, or at least reduced fragmentation of a massive gas cloud, resulting in the formation of a supermassive star (SMS). Both of these mechanisms create a heavy seed with masses in excess of approximately $10^{3}$ $\rm{M_{\odot}}~{}$and in some cases possibly up to $10^{5}$ $\rm{M_{\odot}}$. We will address these questions and gauge their possible number density in §3. We begin by first briefly reviewing the light seed model and then the heavy seed model. More in-depth reviews of the mechanisms can be found in reviews by Latif & Ferrara (2016); Woods et al. (2017); Inayoshi, Visbal & Haiman (2020); Volonteri, Habouzit & Colpi (2021). Additional mechanisms, which we will not consider in this paper, include supermassive stars fuelled by dark matter annihilation (e.g. Spolyar, Freese & Gondolo, 2008; Feng, Yu & Zhong, 2021; Singh, Monaco & Tan, 2023) or primordial black holes (e.g., Carr & Silk, 2018; Hasinger, 2020). We do not consider these mechanisms here as there is still considerable uncertainty in their formation pathway and their possible number densities. However, continued research into these mechanisms is warranted given the outstanding questions over the mainstream pathways, as we now discuss, as well as the ability of "non-standard" dark matter to potentially solve other dynamical hurdles (e.g Alonso-Álvarez, Cline & Dewar, 2024).
@@ 2 Light Seeds as the origin of all MBHs
where $\rm{t_{Edd}}=0.45$ Gyr and $\epsilon_{r}$ is the radiative efficiency. For a “standard” radiative efficiency of $\epsilon_{r}\sim 0.1$ and a black hole seed mass of $\rm{M_{0}}=10^{2}$ $\rm{M_{\odot}}$, it takes nearly 1 Gyr to grow from $\rm{M(0)}$ up to $\sim 10^{9}$ $\rm{M_{\odot}}$. While therefore in principle marginally allowable with current observational constraints the growth must continue at the Eddington rate all the way up to the maximum mass (i.e. the duty cycle must be unity). Such a scenario has been shown to be exceedingly unlikely with numerous models and cosmological simulations showing that light seeds do not grow efficiently and tend to remain at their seed masses across cosmic time (Alvarez, Wise & Abel, 2009; Milosavljević, Couch & Bromm, 2009; Smith et al., 2018; Spinoso et al., 2023).
@@ 3 Heavy Seeds as a Sub-population
We note here that high resolution modelling including state-of-the-art physical models is crucial here. When attempting to model the collapse of gas creating the central MBH. resolution and/or physical models can result in fragmenting objects which are unrealistically massive. Early simulations of PopIII formation in particular suffered from this issue (e.g. Bromm, Coppi & Larson, 1999; Abel, Bryan & Norman, 2000) while more recent approaches (when modelling heavy seed formation) have shown, for example, the importance of resolving dynamical heating (e.g. Wise et al., 2019) and ISM turbulence (e.g. Latif et al., 2022).
@@ 4 Seed Number Densities
Dijkstra, Ferrara & Mesinger (2014) calculated the number density of heavy seeds forming through the so-called Lyman-Werner (LW) channel. In this case an atomic halo is irradiated with a super-critical LW by a nearby galaxy (Dijkstra et al., 2008; Agarwal et al., 2012, 2013; Visbal, Haiman & Bryan, 2014; Agarwal et al., 2016; Regan et al., 2017). In principle this can lead to the formation of an extremely massive seed (see §5). However, the calculated number densities are small with expected number densities in the range $10^{-10}$ cMpc-3 to $10^{-5}$ cMpc-3 (at $z\sim 10$). It should be noted here that the higher number densities quoted here (i.e. $10^{-5}$ cMpc-3) should be treated as strong upper limits as they rely on very optimistic formation criteria, namely, low values of the LW radiation needed to keep $\rm{H_{2}}$ dissociated or allowing for metal pollution. Other authors have drawn similar conclusions on the expected number densities from this channel (Inayoshi & Tanaka, 2015; Habouzit et al., 2016; Agarwal et al., 2016; Trinca et al., 2022).
Moving now to channels which potentially produce heavy seeds below the theoretical maximum mass (of SMSs) but well in excess of light seeds masses. The so-called rapid assembly channel (Yoshida et al., 2003; Fernandez et al., 2014; Wise et al., 2019; Regan et al., 2020c; Lupi, Haiman & Volonteri, 2021; Latif et al., 2022) can produce heavy seeds with number densities in the range 10-5 to $10^{-1}$ cMpc-3 (at $z\sim 10$) (McCaffrey et al. in prep) depending on environmental factors. Number densities achieved through other channels like baryonic streaming velocities result in number densities close to approximately 10-5 cMpc-3 (at $z\sim 10$) as well (Hirano et al., 2017).
@@ 5.1 The 1000J Halo
The number density of haloes subjected to a super-critical LW flux cannot be obtained from the simulations discussed above due to their idealistic nature. However, both Visbal, Haiman & Bryan (2014) and Regan et al. (2020b) (using the Renaissance simulation suite) computed the abundance of so-called synchronised pairs which can lead to the emergence of haloes similar to the 1000J halo (Regan et al., 2017). Both authors found the abundance of synchronised haloes to be less than $10^{-2}$ cMpc-3 (at z $\sim 10$). However, this number should be seen as an upper limit. For the "Normal" region of the Renaissance simulation suite Regan et al. (2020b) found exactly zero synchronised pairs while finding five, in the clustered, "Rarepeak" region. Taking into account the relative rarity of the Rarepeak region, the number densities obtained should be decreased by a factor of approximately 1000. This leads to number densities of less than $10^{-5}$ cMpc-3 (at z $\sim 10$) for a 1000J like object. An important additional point here is that in-situ LW radiation can potentially ease the thresholds somewhat but remains in delicate balance with other factors like metal-pollution which can suppress heavy seed formation (e.g. Dunn et al., 2018; Bhowmick et al., 2022b; Chiaki et al., 2023).
@@ 6 Model Halo Comparison
There are two primary reasons for the elevated stellar masses found in the 1000J compared to the Rapid and PopIII halo. On the one hand and as described above the LW flux dissociates $\rm{H_{2}}$ and leads to larger Jeans masses and hence larger protostars (Prole et al., 2024) but also these protostars can accrete more readily and for longer (see Figure 5). In the Rapid halo and PopIII halo the structures are far more asymmetric making continued accretion much more difficult (at least for the stellar masses probed here). For the cases of stars in both the Rapid and PopIII haloes the stars can initially accrete gas readily but the gas supply quickly gets consumed and the stars contract to the main sequence (Regan et al., 2020c) with masses significantly below that seen in more symmetric and homogeneous halos.
It should be noted that in Figure 5 we show only accretion over a relatively short period (hundreds of kilo-years). The time-resolution needed for these simulations is very short - of the order of years compared to large scale cosmological simulations with timesteps many orders of magnitude higher. The subsequent growth of the seed (light or heavy) will depend fundamentally on the evolution of the host galaxy over several tens and hundreds of megayears. Presently, simulations able to track the model the formation of light and heavy seeds and to subsequently follow their growth are extremely computationally challenging and essentially intractable. We nonetheless describe progress in this direction in the following section bearing in mind that this remains an extremely challenging modelling problem.
Figure 5: The evolution of the most massive star that forms in each of the three haloes (1000J, Rapid and PopIII). Stars are modelled as accreting sink particles. The spatial resolution of each simulation is approximately $10^{-3}$ pc (physical). In the top panel we show the mass as a function of time. In the bottom panel we show the accretion rate onto the most massive star in each halo as a function of time. The 1000J halo forms the most massive object followed by the Rapid halo and the PopIII halo. The dashed line shows the critical accretion that must be maintained onto the stellar surface to meet the criteria for being a super-massive star (Nandal et al., 2023).
@@ 7.1 Growth of Light Seeds
Smith et al. (2018) et al. used the Renaissance suite to track, in post-processing, the evolution of light seed black holes formed from PopIII stars. They tracked the evolution of over 20,000 light seeds for up to 200 Myrs and found negligible growth across the entire light seed population. The reason being is that the light seeds experience extremely low duty cycles due to the surrounding (negative) feedback processes of star formation and stellar feedback. Moreover, higher resolution studies of (PopIII) star forming environments show that PopIII stars form in multiples (e.g. Turk, Abel & O’Shea, 2009; Clark et al., 2011; Prole et al., 2024) setting off a competitive accretion scenario further mitigating against efficient accretion. Other authors (e.g. Pfister et al., 2019; Ma et al., 2021) find that only seeds with masses in the heavy seed (up to $10^{8}$ $\rm{M_{\odot}}~{}$in the case of the Ma et al. (2021) study) range can effectively sink to the centre of the potential well where efficient accretion may be sustainable. None of the above studies rule out light seed accretion but instead find that the probability of efficient light seed growth is extremely low.
Building on these challenging conclusions from studies of light seed growth a number of groups have investigated how growth can be accelerated when the conditions are favourable. In the majority of these cases the light seed black hole is modelled to either sit at the centre of a converging flow (e.g. Volonteri & Rees, 2005) or is "trapped" within the confines of a nuclear star cluster (Alexander & Natarajan, 2014; Natarajan, 2020) or within a circumnuclear disk (e.g. Lupi et al., 2016). More recent work by Shi, Kremer & Hopkins (2024) shows that a small fraction ($\sim$ 1%) of light seeds can achieve bursts of rapid growth when placed in environments that are favourable (i.e. are gas rich and support growth). In each of these cases the growth of the initially light seed proceeds via bursts of super-Eddington accretion. The physical process of super-Eddington accretion has been demonstrated extensively via extremely high resolution simulations (e.g. Sa̧dowski, 2009; Sa̧dowski & Narayan, 2016; Sa̧dowski et al., 2016; Jiang, Stone & Davis, 2019; Jiang et al., 2020). However, what is lacking are robust numerical simulations which show self-consistently the growth of light seeds in a cosmological setting. The reason for this is that a combination of high resolution, with resolved dynamics, and a realisation containing a rare dynamical occurrence is required. This has, thus far, not been realised. However, such a setup should be tractable with current simulation codes and techniques at least for the early evolutionary phase.
@@ 7.2 Heavy Seed Growth
So overall where does the extensive work on black hole growth leave us? Across the various simulation setups there is convergence that black holes can grow once the host halo becomes massive enough with some differences arising for black hole growth in lower mass galaxies which needs to be fully understood. Additionally, the black hole must achieve a mass of at least $10^{5}$ $\rm{M_{\odot}}~{}$before growth can be achieved - this is particularly evident from the Astrid simulations which have the lowest seed masses ($\rm{M_{seed}\gtrsim 10^{4}}$ $\rm{M_{\odot}}$) and see very limited growth for the lightest seeds. Similar results are obtained in semi-analytical models (Trinca et al., 2022; Spinoso et al., 2023). The main result therefore is that any MBH seeds below approximately $10^{5}$ $\rm{M_{\odot}}~{}$grow extremely inefficiently due to dynamics. In addition to this, in low mass galaxies ($\rm{M_{*}\lesssim 10^{9}}$ $\rm{M_{\odot}}~{}$) growth may by hampered by supernova feedback but this result suffers from different interpretations in the literature and requires further study.
@@ 8 Summary and Conclusions
The goal of this paper is to weave together recent developments in modeling MBH seeding and pathways to their growth. The current paradigm is broken into light and heavy seeds with light seeds typically defined as those with initial masses less than 1000 $\rm{M_{\odot}}~{}$and heavy seeds as those with initial masses in excess of 1000 $\rm{M_{\odot}}$(e.g. Sassano et al., 2021; Evans, Blecha & Bhowmick, 2023; Jeon et al., 2024). Scenarios considered in the past gave rise to a bimodal initial mass function: common light seeds related to PopIII stars with mass $\sim 100$ $\rm{M_{\odot}}~{}$and rarer heavy seeds related to SMSs with mass $\sim 10^{5}$ $\rm{M_{\odot}}$. In order to grow these seeds into the massive and supermassive regime, two main pathways were considered: super-Eddington on light seeds and sub-Eddington on heavy seeds, although of course super-Eddington accretion can occur, and indeed it may be more likely to occur, for heavy seeds. What the most recent simulations are instead suggesting is that these two regimes are likely to be the extreme cases of a broad distribution of initial properties.
@@
    Light Seeds: Smaller haloes or those enriched with metals will produce relatively less massive stars. These stars will be distributed inhomogeneously around the halo centre and will accrete very inefficiently. The largest stellar masses shown here have masses up to a few times $\sim 10^{2}$ $\rm{M_{\odot}}~{}$(PopIII halo). The number densities of these halos are between $10^{-1}$ cMpc-3 and up to $10$ cMpc-3 (z $\sim$ 10).
@@
    Need for gas replenishment: In all three halos considered the accretion rate on the stellar progenitors of the seeds drops after $\sim 10^{5}$ years as the readily accretable gas in the parent halo is consumed. Gas replenishment is needed (Johnson & Bromm, 2007; Lupi et al., 2024) through either mergers with gas-rich halos or accretion from the cosmic web.
@@
    Heavy Seed Accretion: Heavy seeds experience the same bottlenecks as light seeds, but their effect is mitigated in relation to the seed masses: Bondi accretion scales as $M_{BH}^{2}$ and gravity scales as $M_{BH}$. A 10-times more massive seed has better chances of remaining in dense regions and accreting from them (Pfister et al., 2019; Lupi et al., 2024).
@@
Figure 6: Schematic comparing Semi Analytic Models (SAMs) against hydrodynamic simulations. The computational complexity increases from the top right to the bottom left. Models in the top right run extremely quickly but can lack the physical complexity to resolve small scale features like MBH dynamics and seed formation. On the other hand models occupying the bottom left panel focus on scales close to the event horizon and can only be evolved for relatively short physical times given the extreme computational expenses. An array of simulations volumes and resolutions must therefore be used to "join" the scales. The simulation images shown include Illustris (e.g. Vogelsberger et al., 2014), Romulus (e.g. Tremmel et al., 2017), Renaissance (e.g. Xu, Wise & Norman, 2013) and results from small scale GRMHD simulations (Chatterjee et al., 2019).